OpenStudy (anonymous):

Write in standard form an equation that has x-intercept of 5 and a y-intercept of -4

OpenStudy (dtan5457):

You know that two of your points must be 5,0 and 0,-4 Use the slope formula.

OpenStudy (dtan5457):

\[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (dtan5457):

-4-0/0-5

OpenStudy (dtan5457):

Your slope would be\[\frac{ 4 }{ 5 }\]

OpenStudy (dtan5457):

Now, knowing your y intercept in slope. Put all that info into y=mx+b